# ISDrama: Immersive Spatial Drama Generation through Multimodal Prompting

### v1

@@ A.3 Mamba
$\mathbf{\overline{A}}$ $=\exp{(\mathbf{\Delta}\mathbf{A})},$ (14)
@@
$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$ (15)
@@
$\mathbf{\overline{K}}$ $=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{A}}\mathbf{\overline{B}},\dots,\mathbf{C}\mathbf{\overline{A}}^{\mathtt{M}-1}\mathbf{\overline{B}}),$ (16)
@@ Appendix B Dataset Details
Figure 5: The extended statistics of MRSDrama.
Figure 6: The binaural recording device we used.
Simulated data cannot accurately capture the intricate, dramatic prosody or the precise effects of real-world spatial scenes, positions, and orientations. Meanwhile, stimulated spatial audio datasets that expand on monaural audio data also have limited effectiveness in modeling the nuances of real-world scenes, positions, and orientations (Sarabia et al., 2023; Li et al., 2024a).
@@
Annotators label the 3D coordinates and quaternion orientations of the sound source (the speaker’s mouth) based on the video. Using the annotated pose data, we employ GPT-4o (Achiam et al., 2023) to generate textual prompts, such as: ”Speaker 1 initially stands in the front-right, then walks at a slow pace to the far front-left…” For binaural speech, we perform denoising using FRCRN (Zhao et al., 2022) and extract phonemes using PyPinyin. Following previous Chinese audio annotation works (Zhang et al., 2024a, b, c; Li et al., 2024b; Guo et al., 2025; Zhang et al., 2022). Coarse alignment between phonemes and audio is achieved using MFA (McAuliffe et al., 2017). Annotators then refine the rough alignment using Praat (Boersma, 2001), focusing on correcting word and phoneme boundaries and addressing erroneous words based on auditory perception. Each step is double-checked by human annotators to ensure accuracy. We hire annotators at a rate of \$30 per hour, and they consent to their contributions being used for research purposes. Finally, the dataset is segmented into 47,958 segments according to speaker transitions, with each segment having a maximum duration of 16 seconds. Figure 5 shows more statistics of MRSDrama on movement speeds and segment duration.
@@ C.1 Model Configuration
For the spatial audio encoder and decoder, we use a Variational Autoencoder (VAE) architecture (Kingma & Welling, 2013). Mel-spectrograms are derived from binaural waveforms with a 48 kHz sample rate, 1024 window size, 256 hop size, and 80 mel bins. We use HiFi-GAN (Kong et al., 2020) as the vocoder to synthesize waveforms from mel-spectrograms. Specifically, the model consists of 3 layers for both the encoder and decoder, with a hidden size of 384 and a Conv1D kernel size of 5. The binaural mel-spectrogram with dimensions $B,2,80,T$ is compressed into dimensions $B,40,T/4$, facilitating further processing by the Transformer. During training, batches of fixed length are used, consisting of 3000 mel-spectrogram frames. The Adam optimizer is employed with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and 10K warm-up steps.
@@ C.3 Inference Procedure
C.4 Multimodal Pose Encoder
@@
C.5 Spatial Audio Encoder and Decoder
The Spatial Audio Encoder and Decoder are designed based on the Variational Autoencoder (VAE) model (Kingma & Welling, 2013). During pre-training, we optimize the encoder and decoder using the L2 reconstruction loss. To further enhance reconstruction quality, we integrate a GAN discriminator inspired by the architecture of ML-GAN (Chen et al., 2020). Specifically, we employ the LSGAN-style adversarial loss (Mao et al., 2017), $\mathcal{L}_{adv}$, which minimizes the distributional divergence between the predicted mel-spectrograms and the ground truth mel-spectrograms. Before encoding, we extract the mel-spectrogram using librosa  [^2] https://github.com/librosa/librosa. After generating the mel-spectrogram from the decoder’s output, HiFi-GAN (Kong et al., 2020) is used to convert it back into an audio waveform. To improve the quality of speech generation, we add the reconstructed F0 from our Immersive Drama Transformer during inference, applying the neural source filter (NSF) strategy for enhanced quality.
C.6 Mamba-Transformer Block
@@
To enhance training stability and prevent numerical instability caused by uncontrolled growth in absolute values, we adopt RMSNorm (Zhang & Sennrich, 2019). For encoding scene $s$, we encode a video frame or textual scene descriptions into scene embedding $z_{s}$. Subsequently, the global embedding $z_{g}$ is computed by averaging the prompt audio embedding $z_{a}$ and the scene embedding $z_{s}$ along the temporal dimension, followed by the addition of the time step embedding $z_{t}$. This global embedding is processed through a global adaptor, which modulates the latent representation via adaptive layer normalization (AdaLN) (Peebles & Xie, 2023) to ensure style consistency. The scale and shift parameters are calculated using linear regression:
$AdaLN(h,c)=\gamma_{c}\times LayerNorm(h)+\beta_{c},$ (17)
where $h$ represents the hidden representation. The batch normalization scale factor $\gamma$ is zero-initialized in each block (Peebles & Xie, 2023). Additionally, we utilize rotary positional embeddings (RoPE) (Su et al., 2024) as a form of relative positional encoding, injecting temporal positional information into the model. This enhances the model’s ability to capture temporal relationships between sequential frames, leading to notable performance gains in the transformer.
@@
$Attention=\text{softmax}\left(\frac{\tilde{Q_{h}}\tilde{K_{h}}^{\top}}{\sqrt{d}}\right)V_{h}+$ (18)
@@ C.7 Drama-MOE
$g(h)_{i}=\frac{\exp((h\cdot W_{g}+\zeta_{i})/\tau)}{\sum_{j=1}^{N}\exp((h\cdot W_{g}+\zeta_{j})/\tau)},$ (19)
@@
Figure 7: Screenshot of monaural speech evaluation.
Figure 8: Screenshot of binaural speech evaluation.
@@ Appendix D Evaluation Metrics
D.1 Objective Evaluation
@@
$S_{i}(t,m)=\log\left(|X_{i}(t,f)|^{2}\times\text{melW}\right),$ (21)
@@
$ILD(t,f)=20\log_{10}\left(\dfrac{|X_{2}(t,f)|+\varepsilon}{|X_{1}(t,f)|+\varepsilon}\right),\varepsilon=1e^{-10}.$ (23)
@@
Figure 9: Visualization results.
D.2 Subjective Evaluation
@@ E.1 Baseline Models
StyleTTS 2 (Li et al., 2024c) combines style diffusion and adversarial training with large speech language models (SLMs) for high-quality text-to-speech synthesis. It models style as a latent random variable using diffusion models. We employ their official code  [^5] https://github.com/yl4579/StyleTTS2.
@@ E.2 Visualization Results
E.3 Drama-MOE
Figure 10: The statistics of spatial routing in Drama-MOE.
Table 5: Ablation study for Drama-MOE.
$\bf Expert$ CMOS-Q CMOS-E
1 -0.63 -0.79
2 -0.47 -0.44
3 -0.19 -0.25
4 0.00 0.00
5 -0.02 0.07
For the Spatial-MOE component in Drama-MOE, we find that expert selection is highly influenced by the horizontal angle (azimuth). To analyze this, we evaluate the softmax output probabilities before expert selection.
As shown in Figure 10, we only present results for the 0°–180° azimuth range, as expert selection patterns for the front (0°–180°) and back (180°–360°) azimuths are highly similar. Thus, we aggregate them for clarity. Our results reveal a distinct expert specialization based on azimuth: 1) Expert 1 primarily covers 0° to 45°. 2) Expert 2 is most active in the 45° to 120° range. 3) Expert 3 focuses on 1200° to 135°. 4) Expert 4 is dominant in 135° to 180°. This distribution demonstrates the effectiveness of our load-balancing strategy, ensuring that Spatial-MOE effectively selects different experts based on spatial information.
Furthermore, we conduct an ablation study on the number of experts in Spatial-MOE, as shown in Table 5. To evaluate perceptual differences, we performed a CMOS (Comparative Mean Opinion Score) test.
Our results indicate a consistent improvement in generation quality as the number of experts increases. However, beyond four experts, the gains become marginal or negligible. This is likely due to two factors: 1) Increased model complexity and learning burden. An excessive number of experts may introduce unnecessary redundancy, making it harder for the model to converge effectively. 2) Higher computational cost. More experts lead to significantly increased training and inference overhead without proportional gains in performance. Based on these findings, we set the number of experts per group to four, achieving a balance between performance and efficiency.
E.4 Context-Consistent CFG
Following previous works about CFG (Jiang et al., 2025), we conduct experiments to verify the parameter settings of Equation 5, as shown in Table 6. For evaluation, we implement CMOS assessments. When $\alpha=1$, $v_{cfg}$ relies solely on the input prompt audio. Moreover, if $\alpha=1$ and $\gamma=1$, $v_{cfg}$ becomes equivalent to the original formulation $v_{t}(x,t|a,C;\theta)$.
Table 6: Ablation study for context-consistent CFG.

### v6

@@ A.3. Mamba
(14) $\mathbf{\overline{A}}$ $=\exp{(\mathbf{\Delta}\mathbf{A})},$
@@
(15) $h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$
@@
(16) $\mathbf{\overline{K}}$ $=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{A}}\mathbf{\overline{B}},\dots,\mathbf{C}\mathbf{\overline{A}}^{\mathtt{M}-1}\mathbf{\overline{B}}),$
@@ Appendix B Dataset Details
Figure 5. The extended statistics of MRSDrama.
[…]
The extended statistics of MRSDrama.
Figure 6. The binaural recording device we used.
[…]
The binaural recording device we used.
Simulated data cannot accurately capture the intricate, dramatic prosody or the precise effects of real-world spatial scenes, positions, and orientations. Meanwhile, stimulated spatial audio datasets that expand on monaural audio data also have limited effectiveness in modeling the nuances of real-world scenes, positions, and orientations (Sarabia et al., 2023; Li et al., 2024b).
@@
Annotators label the 3D coordinates and quaternion orientations of the sound source (the speaker’s mouth) based on the video. Using the annotated pose data, we employ GPT-4o (Achiam et al., 2023) to generate textual prompts, such as: ”Speaker 1 initially stands in the front-right, then walks at a slow pace to the far front-left…” For binaural speech, we perform denoising using FRCRN (Zhao et al., 2022) and extract phonemes using PyPinyin. Following previous Chinese audio annotation works (Zhang et al., 2024a, b, c; Li et al., 2024c; Guo et al., 2025; Zhang et al., 2022). Coarse alignment between phonemes and audio is achieved using MFA (McAuliffe et al., 2017). Annotators then refine the rough alignment using Praat (Boersma, 2001), focusing on correcting word and phoneme boundaries and addressing erroneous words based on auditory perception. Each step is double-checked by human annotators to ensure accuracy. We hire annotators at a rate of \$30 per hour, and they consent to their contributions being used for research purposes. Finally, the dataset is segmented into 47,958 segments according to speaker transitions, with each segment having a maximum duration of 16 seconds. Figure 5 shows more statistics of MRSDrama on movement speeds and segment duration.
@@ C.1. Model Configuration
For the spatial audio encoder and decoder, we use a Variational Autoencoder (VAE) architecture (Kingma and Welling, 2013). Mel-spectrograms are derived from binaural waveforms with a 48 kHz sample rate, 1024 window size, 256 hop size, and 80 mel bins. We use HiFi-GAN (Kong et al., 2020) as the vocoder to synthesize waveforms from mel-spectrograms. Specifically, the model consists of 3 layers for both the encoder and decoder, with a hidden size of 384 and a Conv1D kernel size of 5. The binaural mel-spectrogram with dimensions $B,2,80,T$ is compressed into dimensions $B,40,T/4$, facilitating further processing by the Transformer. During training, batches of fixed length are used, consisting of 3000 mel-spectrogram frames. The Adam optimizer is employed with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and 10K warm-up steps.
@@ C.3. Inference Procedure
C.4. Multimodal Pose Encoder
@@
C.5. Spatial Audio Encoder and Decoder
The Spatial Audio Encoder and Decoder are designed based on the Variational Autoencoder (VAE) model (Kingma and Welling, 2013). During pre-training, we optimize the encoder and decoder using the L2 reconstruction loss. To further enhance reconstruction quality, we integrate a GAN discriminator inspired by the architecture of ML-GAN (Chen et al., 2020). Specifically, we employ the LSGAN-style adversarial loss (Mao et al., 2017), $\mathcal{L}_{adv}$, which minimizes the distributional divergence between the predicted mel-spectrograms and the ground truth mel-spectrograms. Before encoding, we extract the mel-spectrogram using librosa  [^2] https://github.com/librosa/librosa. After generating the mel-spectrogram from the decoder’s output, HiFi-GAN (Kong et al., 2020) is used to convert it back into an audio waveform. To improve the quality of speech generation, we add the reconstructed F0 from our Immersive Drama Transformer during inference, applying the neural source filter (NSF) strategy for enhanced quality.
C.6. Mamba-Transformer Block
@@
To enhance training stability and prevent numerical instability caused by uncontrolled growth in absolute values, we adopt RMSNorm (Zhang and Sennrich, 2019). For encoding scene $s$, we encode a video frame or textual scene descriptions into scene embedding $z_{s}$. Subsequently, the global embedding $z_{g}$ is computed by averaging the prompt audio embedding $z_{a}$ and the scene embedding $z_{s}$ along the temporal dimension, followed by the addition of the time step embedding $z_{t}$. This global embedding is processed through a global adaptor, which modulates the latent representation via adaptive layer normalization (AdaLN) (Peebles and Xie, 2023) to ensure style consistency. The scale and shift parameters are calculated using linear regression:
(17) $AdaLN(h,c)=\gamma_{c}\times LayerNorm(h)+\beta_{c},$
where $h$ represents the hidden representation. The batch normalization scale factor $\gamma$ is zero-initialized in each block (Peebles and Xie, 2023). Additionally, we utilize rotary positional embeddings (RoPE) (Su et al., 2024) as a form of relative positional encoding, injecting temporal positional information into the model. This enhances the model’s ability to capture temporal relationships between sequential frames, leading to notable performance gains in the transformer.
@@
(18) $Attention=\text{softmax}\left(\frac{\tilde{Q_{h}}\tilde{K_{h}}^{\top}}{\sqrt{d}}\right)V_{h}+$
@@ C.7. Drama-MOE
(19) $g(h)_{i}=\frac{\exp((h\cdot W_{g}+\zeta_{i})/\tau)}{\sum_{j=1}^{N}\exp((h\cdot W_{g}+\zeta_{j})/\tau)},$
@@ Appendix D Evaluation Metrics
Figure 7. Screenshot of monaural speech evaluation.
[…]
Screenshot of monaural speech evaluation.
Figure 8. Screenshot of binaural speech evaluation.
[…]
Screenshot of binaural speech evaluation.
D.1. Objective Evaluation
@@
(21) $S_{i}(t,m)=\log\left(|X_{i}(t,f)|^{2}\times\text{melW}\right),$
@@
[…]
@@
D.2. Subjective Evaluation
@@ E.1. Baseline Models
StyleTTS 2 (Li et al., 2024a) combines style diffusion and adversarial training with large speech language models (SLMs) for high-quality text-to-speech synthesis. It models style as a latent random variable using diffusion models. We employ their official code  [^5] https://github.com/yl4579/StyleTTS2.
@@ E.2. Visualization Results
E.3. Context-Consistent CFG
We conduct experiments to verify the parameter settings of Equation 5, as shown in Table 5. For evaluation, we implement CMOS assessments. When $\alpha=1$, $v_{cfg}$ relies solely on the input prompt audio. Moreover, if $\alpha=1$ and $\gamma=1$, $v_{cfg}$ becomes equivalent to the original formulation $v_{t}(x,t|a,C;\theta)$.
Table 5. Ablation study for context-consistent CFG.